○소 Physics II SP2019 H01 1 of1 n Iron Rod On a part-time job, you are asked to bring a cylindrical iron rod of density 7800 kg/mlength 81.1 cm and dameter acceleration due to gravity, g 9.81 m/ 2.40 em from a storage room to a machinist Calculate the weight of the rod, uThe Express your answer numerically in newtons Part B Next >

On a part-time job, you are asked to bring a cylindrical iron rod of density 7800 kg length 81.1 cm and diameter acceleration due to gravity, g 9.81 m.
Answer #1

area of rod = (pi) (.024/2)^2*0.811 = 0.0003669 m^2

mass = area * density

= 0.0003669 * 7800= 2.86 kg

weight = mass*gravity

= 2.86 kg* 9.81 m/s^2

= 28.07 N